# Unusual intra-burst variations of polarization states in FRB 20210912A and FRB 20230708A : Effects of plasma birefringence?

### v1

@@ Abstract
Fast radio bursts (FRBs) are highly energetic events of short-duration intense radio emission, the origin of which remains elusive till date. Polarization of the FRB signals carry information about the emission source as well as the magneto-ionic media the signal passes through before reaching terrestrial radio telescopes. Currently known FRBs show a diverse range of polarization, sometimes with complex features, making it challenging to describe them in a unified model. FRB 20230708A and FRB 20210912A are two bright and highly polarized FRBs detected in the Commensal Real-time ASKAP Fast Transients (CRAFT) survey with the Australian Square Kilometre Array Pathfinder (ASKAP) that exhibit time-dependent conversion between linear and circular polarizations as well as intra-burst (apparent) variation of Faraday rotation measure. We investigate the intra-burst temporal evolution of the polarization state of radio emission in these two events using the Poincaré sphere representation and find that the trajectories of the polarization state are well described by great circles on the Poincaré sphere. These polarization features may be signatures of a transition between two partially coherent orthogonal polarization modes or propagation through a birefringent medium. We find that the observed variations of the polarization states of these two FRBs are qualitatively consistent a magnetospheric origin of the bursts and the effects of propagation through a birefringent medium with linearly polarized modes in the outer magnetosphere or near-wind region of a neutron star.
@@ 1 Introduction
A large fraction ($\gtrsim 60\%$) of the known FRBs are highly polarized; while the vast majority show predominantly linear polarization, a significant degree of circular polarization has been observed in $\sim$20of FRBs (e.g. Sherman et al., 2023). Some FRBs exhibit intra-burst temporal variation of the polarization state of radio emission (e.g. Cho et al., 2020; Sherman et al., 2023); the reason behind this is still not well understood. In this work, we investigate the intra-burst short timescale ($\sim$ 10 $\mu$s) variation of polarization state in two ASKAP-detected FRBs — FRB 20230708A (Dial et al., submitted; Shannon et al., 2024) and FRB 20210912A (Marnoch et al., 2023; Bera et al., 2024) — which show significant and varying degrees of linear and circular polarizations, as well as “apparent” variation of Faraday Rotation measure [^1] Throughout this work we use “apparent RM”/“observed RM” to refer to the slope of the observed position angle ($\rm PA_{obs}$) vs $\lambda^{2}$ curve. This quantity may not always be physically associated with Faraday rotation – the effects of propagation through a cold, magnetized, electron-ion plasma, which has circularly polarized propagation modes., across their burst profiles. The Poincaré sphere representation provides a convenient way for analysis and interpretation of such variations of the polarization state of electromagnetic waves (e.g. Radhakrishnan, 1994; Malykin, 1997). We use the Poincaré sphere representation to analyze the temporal evolution of the polarization vectors of FRB 20230708A and FRB 20210912A, and discuss possible interpretations for the observed variations of the polarization states.
@@ 2 FRB Data and Analysis Methods
FRB 20230708A (Dial et al., submitted) and FRB 20210912A (Marnoch et al., 2023; Bera et al., 2024) were detected in the Commensal Real-time ASKAP Fast Transients (CRAFT; Bannister et al., 2017; Shannon et al., 2024) survey on the Australian Square Kilometre Array Pathfinder (ASKAP; Hotan et al., 2021), using the Fast Real-time Engine for Dedispersing Amplitudes (FREDDA; Qiu et al., 2023). The raw voltage data recorded upon detection were processed offline using the CRAFT Effortless Localization and Enhanced Burst Inspection pipeline (CELEBI; Scott et al., 2023). Important observed and inferred properties of these two FRBs are listed in Table 1. Polarization calibration was done as part of post-processing, using the Vela pulsar (PSR J0835-4510) as the calibrator, for both FRBs. Details of the polarization analysis — including construction of the dynamic spectra of Stokes’ parameters I, Q, U, V and calculation of the linearly polarized flux density ($L$) — are described in Bera et al. (2024) and Dial et al. (submitted). The observed position angle (PA, $\chi_{\rm obs}$) of linear polarization, and the polarization ellipticity angle (EA, $\psi_{\rm obs}$), were calculated using the relations
@@ 2.1 Faraday Rotation
Figure 1: Frequency averaged polarization time profile of FRB 20230708A in the primary sub-burst [A, left] and in the fainter trailing sub-bursts [B, right] after correcting for $\rm RM=-5.75\;rad\>m^{2}$. The top panels show the normalized profiles of total intensity (I), linear polarization (L) and circular polarization (V), as well as estimates of RM in different time ranges (indicated by the x-errorbars). The middle panels show profiles of fractional total (P), linear (L) and circular (V) polarization. The bottom panels show polarization position angle (PA, $\chi$) and ellipticity angle (EA, $\psi$). The time axis, with its reference (zero) set to the maximum of the intensity profile, has different scales in the left and the right panels.
@@ 2.2 Polarization Time Profiles
We used the Poincaré sphere representation (e.g. Malykin, 1997), which is a unit sphere in the 3-dimensional polarization space with (Cartesian) co-ordinate axes defined by $Q/I$, $U/I$ and $V/I$, to investigate the temporal changes in the polarization state of the FRBs. Each point on the surface of the Poincaré sphere represents a unique polarization state of fully polarized electromagnetic (EM) waves, with longitude $=2\chi$ and latitude $=2\psi$. Diametrically opposite points represent mutually orthogonal polarization states. Partially polarized EM waves can be represented by points beneath the surface, i.e. with radii smaller than unity. All FRBs discussed in this work are highly polarized with total polarization fractions $\gtrsim 70\%$ over their entire burst profiles. We hence neglect the “depth” (beneath the surface) and plot all points on the surface of the Poincaré sphere using the measured values of $\chi$ and $\psi$.
@@ 3 FRB 20230708A
FRB 20230708A is a relatively bright event featuring a double-peaked primary sub-burst followed by trailing faint quasi-periodic sub-bursts, as shown in Figure 1 (see Dial et al., submitted, for further details). The primary sub-burst has an average rotation measure of $\rm RM_{avg}=-5.75\pm 0.20\;rad\>m^{2}$, which is consistent with the RM estimates for the trailing sub-bursts [^2] The same data that has been used by Dial et al. (submitted) were independently analysed for this work. The results are consistent with those presented in Dial et al. (submitted). The difference in the reported values of the average RM resulted from the choice of slightly different time windows for RM estimation and the “apparent” temporal variation of RM.. However, polarization spectra integrated over shorter time ranges show a smooth temporal variation of apparent RM across the primary sub-burst.
@@ 3.1 Trajectory on the Poincaré Sphere
Figure 5: Frequency averaged polarization time profile of FRB 20210912A in sub-burst $A$ [A, left] and sub-burst $B$ [B, right] after correcting for $\rm RM=4.55\;rad\>m^{2}$. The time axis has different scales in the left and the right panels. The vertical dashed line at $t=0.06$ ms in the left panel separates the ‘peak’ and the ‘tail’ of the sub-burst $A$, as discussed in Section 4. See Figure 1 for a description of the symbols.
@@ 4 FRB 20210912A
Figure 6: Temporal evolution of polarization state across the ‘peak’ of sub-burst $A$ of FRB 20210912A on the Poincaré sphere after correcting for $\rm RM=4.55\;rad\>m^{2}$ shown in a 3-D plot (left) and in gnomonic projection (right). Each data point represents the polarization state in a 3.8 $\mu$s time bin. The best-fit great circle is shown as a solid magenta curve in the left panel and a dashed black line in the right panel.
FRB 20210912A is another bright FRB featuring a strong primary sub-burst ($A$) followed by a relatively weaker secondary sub-burst ($B$), as shown in Figure 5 (see Bera et al., 2024, for further details). The average RM of FRB 20210912A, measured over the entire burst profile, is $\rm RM_{avg}=4.55\pm 0.49\;rad\>m^{2}$. However, as discussed in Bera et al. (2024), the apparent RM varies across the two sub-bursts (see the top panels in Figure 5). We de-rotated the $Q-U$ dynamic spectra of FRB 20210912A for $\rm RM_{avg}=4.55\;rad\>m^{2}$ following the method described in Section 2.1. The frequency-averaged time profile shows high degree of polarization and both sub-bursts exhibit time-dependent gradual conversion between linear and circular polarization (middle panels in Figure 5). The PA ($\chi$) and the EA ($\psi$) show smooth correlated temporal evolution across both sub-bursts, although the nature of correlation is different in the two sub-bursts (bottom panels in Figure 5).
As noted by Bera et al. (2024), the fading ‘tail’ of sub-burst $A$ consists of a trailing component with different spectral and polarization properties compared to the ‘peak’ of the sub-burst. The PA and EA trends can be seen to change after the trailing component becomes dominant in the time profile. We hence investigate the polarization properties of the ‘peak’ and ‘tail’ of sub-burst $A$ separately, using a boundary at $t=0.06$ ms (with respect to the maximum of the total intensity profile) that roughly separates the trailing component from the brightest component. We note that sub-burst $B$ also exhibits changes in PA and EA trends at the tail; however, its relative faintness does not allow us to identify any clear component separation.
@@ 4.1 Polarization State in the Two Sub-bursts
We used a time resolution of 3.8 $\mu$s to probe the temporal evolution of polarization state in the peak of sub-burst $A$ of FRB 20210912A. Like the primary sub-burst of FRB 20230708A, the frequency-averaged trajectory of the polarization state of the peak of sub-burst $A$, plotted on the surface of the Poincaré sphere following the method described in Section 2.2, is found to be well described by a great circle, as shown in Figure 6. The best-fit great circle has an inclination of $\Delta_{0}^{A}=62.4\pm 1.4^{\circ}$ and a reference PA of $\chi_{0}^{A}=-10.4\pm 1.3^{\circ}$.
@@
Figure 7: Temporal evolution of polarization state across the ‘tail’ of sub-burst $A$ [A, left] and across sub-burst $B$ [B, right] of FRB 20210912A on the Poincaré sphere after correcting for $\rm RM=4.55\;rad\>m^{2}$. Each data point represents the polarization state of emission in a 7.6 $\mu$s time bin in the left panel and 46 $\mu$s time bin in the right panel. The best-fit great circle is shown as a dashed black line in each panel.
@@ 5 Possible Interpretations
The observed trajectories of polarization state in FRB 20230708A and FRB 20210912A — which results from correlated variations of linear and circular polarization — are unlikely to be a random coincidence, although a formal estimation of the probability is not straightforward and critically depends on the assumptions made. Instrumental effects also appear unlikely to lead to the observed trajectories of the polarization vector as discussed in detail in Section 7.1.
The origin of the circular polarization in FRBs — whether it is an intrinsic property of radio emission or an effect of propagation through magneto-ionic media (e.g. McKinnon, 2024; Kennett & Melrose, 1998; Melrose & Luo, 2004; Lyutikov, 2022; Macquart & Melrose, 2000a, b) — remains unclear. Continuous variation of the polarization state across the burst, with changing fractions of linear and circular polarization and the polarization vector tracing smooth curves on the Poincaré sphere, has been previously observed for other FRBs (Sherman et al., 2023; Cho et al., 2020). Mandlik (2024) reported observation of FRB 20221128A with the UTMOST-NS facility at the Molonglo radio telescope, which shows the polarization vector tracing a segment of a great circle on the Poincaré sphere as seen in FRB 20230708A and FRB 20210912A. Circular polarization induced by scintillation in a turbulent magnetized medium (Macquart & Melrose, 2000b) is unlikely to lead to such smooth trajectories on the Poincaré sphere. ‘Generalized’ Faraday rotation (or Faraday conversion) has been suggested as a possible origin of varying degree of circular polarization (Cho et al., 2020; Kumar et al., 2022; Uttarkar et al., 2024). However, as discussed by Dial et al. (submitted), the time-averaged polarization spectra of FRB 20230708A do not show any conclusive evidence of ‘generalized’ Faraday rotation (Lower, 2021; Uttarkar et al., 2024). Here we discuss three possible scenarios that may lead to the observed great circle trajectories of the polarization vector on the Poincaré sphere for FRB 20230708A and FRB 20210912A — transition between two orthogonal polarization modes and the effects of magneto-ionic birefringence.
@@ 5.2 Effects of Elliptical Birefringence
FRB progenitors are thought to be surrounded by a magneto-ionic shell/nebula, such as a supernova remnant or pulsar wind nebula (e.g. Chatterjee et al., 2017; Michilli et al., 2018; Bruni et al., 2024). Vedantham & Ravi (2019) and Gruzinov & Levin (2019) have suggested that linearly polarized FRB emission can be (partially) converted to circular polarization as the burst propagates through the nebula. These polarization models generally assume that the plasma in the nebula is comprised of a mixture of thermal and relativistic particles, such that the plasma’s natural modes of wave propagation are elliptically polarized. The polarization ellipticity is determined by the ratio of the number densities of the thermal and relativistic particles in the plasma (Kennett & Melrose, 1998). The ellipticity is $\epsilon=\tan(\psi_{o})$, where $\psi_{o}$ is the ellipticity angle of the mode polarization vectors. If the modes are linearly polarized, $\psi_{o}=0$, $\epsilon=0$, and the plasma particles are relativistic. If the modes are circularly polarized, $\psi_{o}=\pm\pi/4$, $\epsilon=\pm 1$, and the plasma particles are thermal, as in the interstellar medium. An elliptically birefringent medium alters the phase offset, $\eta$, between the propagation modes and generally causes the trajectory of the observed polarization vector to trace a small circle on the Poincaré sphere (see, e.g., Figure 3 of Kennett & Melrose, 1998). The mode polarization vectors form the normal to the plane defined by the small or great circle, and the motion along the circle is associated with temporal variation of the phase offset, $\eta$. The relationship between $\psi_{o}$ and the observed inclination angles of the great circles (GCs), $\Delta$, is therefore $\psi_{o}=(\pi/2-\Delta)/2$. The ellipticity determined from the inclination angles observed in FRB 20230708A and the primary sub-burst of FRB 20210912A is $\epsilon\approx 0.25$, indicating the plasma particles are primarily relativistic.
@@
The observed GCs constrain the geometry of the mode polarization vectors relative to the polarization vector of the radiation incident upon the plasma. For the vector to trace a GC, the polarization vector of the incident radiation must be perpendicular to the mode polarization vectors. If the incident radiation is initially linearly polarized, its polarization vector is further constrained to reside along the line formed by the intersection of the GC with the equatorial plane of the sphere.

### v2

@@ Abstract
Fast radio bursts (FRBs) are highly energetic events of short-duration intense radio emission, the origin of which remains elusive till date. Polarization of the FRB signals carry information about the emission source as well as the magneto-ionic media the signal passes through before reaching terrestrial radio telescopes. Currently known FRBs show a diverse range of polarization, sometimes with complex features, making it challenging to describe them in a unified model. FRB 20230708A and FRB 20210912A are two bright and highly polarized (apparently) one-off FRBs detected in the Commensal Real-time ASKAP Fast Transients (CRAFT) survey with the Australian Square Kilometre Array Pathfinder (ASKAP) that exhibit time-dependent conversion between linear and circular polarizations as well as intra-burst (apparent) variation of Faraday rotation measure. We investigate the intra-burst temporal evolution of the polarization state of radio emission in these two events using the Poincaré sphere representation and find that the trajectories of the polarization state are well described by great circles on the Poincaré sphere. These polarization features may be signatures of a transition between two partially coherent orthogonal polarization modes or propagation through a birefringent medium. We find that the observed variations of the polarization states of these two FRBs are qualitatively consistent with a magnetospheric origin of the bursts and the effects of propagation through a birefringent medium with linearly polarized modes located close to the emission source — likely in the outer magnetosphere or near-wind region of a neutron star.
@@ 1 Introduction
A large fraction ($\gtrsim 60\%$) of the known FRBs are highly polarized; while the vast majority show predominantly linear polarization, a significant degree of circular polarization has been observed in $\sim$20of FRBs (e.g. Sherman et al., 2023). Some FRBs exhibit intra-burst temporal variation of the polarization state of radio emission (e.g. Cho et al., 2020; Sherman et al., 2023); the reason behind this is still not well understood. In this work, we investigate the intra-burst short timescale ($\sim$ 10 $\mu$s) variation of polarization state in two ASKAP-detected (apparently) one-off FRBs — FRB 20230708A (Dial et al., 2025; Shannon et al., 2024) and FRB 20210912A (Marnoch et al., 2023; Bera et al., 2024) — which show significant and varying degrees of linear and circular polarizations, as well as “apparent” variation of Faraday Rotation measure [^1] Throughout this work we use “apparent RM”/“observed RM” to refer to the slope of the observed position angle ($\rm PA_{obs}$) vs $\lambda^{2}$ curve. This quantity may not always be physically associated with Faraday rotation; when the PA has a different dependence on $\lambda$ (e.g. Kennett & Melrose, 1998; Lyutikov, 2022) the apparent RM only represents the local slope of the relation within the observed frequency band., across their burst profiles. The Poincaré sphere representation provides a convenient way for analysis and interpretation of such variations of the polarization state of electromagnetic (EM) waves (e.g. Radhakrishnan, 1994; Malykin, 1997). We use the Poincaré sphere representation to analyze the temporal evolution of the polarization vectors of FRB 20230708A and FRB 20210912A, and discuss possible interpretations for the observed variations of the polarization states.
@@ 2 FRB Data and Analysis Methods
FRB 20230708A (Dial et al., 2025) and FRB 20210912A (Marnoch et al., 2023; Bera et al., 2024) were detected in the Commensal Real-time ASKAP Fast Transients (CRAFT; Bannister et al., 2017; Shannon et al., 2024) survey on the Australian Square Kilometre Array Pathfinder (ASKAP; Hotan et al., 2021), using the Fast Real-time Engine for Dedispersing Amplitudes (FREDDA; Qiu et al., 2023). The raw voltage data recorded upon detection were processed offline using the CRAFT Effortless Localization and Enhanced Burst Inspection pipeline (CELEBI; Scott et al., 2023). Important observed and inferred properties of these two FRBs are listed in Table 1. Polarization calibration was done as part of post-processing, using the Vela pulsar (PSR J0835-4510) as the calibrator, for both FRBs. Details of the polarization analysis — including construction of the dynamic spectra of Stokes’ parameters I, Q, U, V and calculation of the linearly polarized flux density ($L$) — are described in Bera et al. (2024) and Dial et al. (2025). The observed position angle (PA, $\chi_{\rm obs}$) of linear polarization, and the polarization ellipticity angle (EA, $\psi_{\rm obs}$), were calculated using the relations
@@ 2.1 Faraday Rotation
Figure 1: Frequency averaged polarization time profile of FRB 20230708A in the primary sub-burst [A, left] and in the fainter trailing sub-bursts [B, right] after correcting for $\rm RM=-5.75\;rad\>m^{2}$. The top panels show the normalized profiles of total intensity (I), linear polarization (L) and circular polarization (V), as well as estimates of RM in different time ranges (indicated by the x-errorbars). Note that the intensity axis has not been shown explicitly. The middle panels show profiles of fractional total (P), linear (L) and circular (V) polarization. The bottom panels show polarization position angle (PA, $\chi$) and ellipticity angle (EA, $\psi$). The time axis, with its reference (zero) set to the maximum of the intensity profile, has different scales in the left and the right panels.
@@ 2.2 Polarization Time Profiles
We used the Poincaré sphere representation (e.g. Malykin, 1997), which is a unit sphere in the 3-dimensional polarization space with (Cartesian) co-ordinate axes defined by $Q/I$, $U/I$ and $V/I$, to investigate the temporal changes in the polarization state of the FRBs. Each point on the surface of the Poincaré sphere represents a unique polarization state of fully polarized EM waves, with longitude $=2\chi$ and latitude $=2\psi$. Diametrically opposite points represent mutually orthogonal polarization states. Partially polarized EM waves can be represented by points beneath the surface, i.e. with radii smaller than unity. All FRBs discussed in this work are highly polarized with total polarization fractions $\gtrsim 70\%$ over their entire burst profiles. We hence neglect the “depth” (beneath the surface) and plot all points on the surface of the Poincaré sphere using the measured values of $\chi$ and $\psi$.
@@ 3 FRB 20230708A
FRB 20230708A is a relatively bright event featuring a double-peaked primary sub-burst followed by trailing faint quasi-periodic sub-bursts, as shown in Figure 1. Dynamic spectra and other details of this FRB have been presented in Dial et al. (2025). The primary sub-burst has an average rotation measure of $\rm RM_{avg}=-5.75\pm 0.20\;rad\>m^{2}$, which is consistent with the RM estimates for the trailing sub-bursts [^2] The same data that has been used by Dial et al. (2025) were independently analyzed for this work. The results are consistent with those presented in Dial et al. (2025). The difference in the reported values of the average RM resulted from the choice of slightly different time windows for RM estimation and the “apparent” temporal variation of RM.. However, polarization spectra integrated over shorter time ranges show a smooth temporal variation of apparent RM across the primary sub-burst.
@@ 3.1 Trajectory on the Poincaré Sphere
Figure 5: Frequency averaged polarization time profile of FRB 20210912A in sub-burst $A$ [A, left] and sub-burst $B$ [B, right] after correcting for $\rm RM=4.55\;rad\>m^{2}$. Note that the intensity axis has not been explicitly shown in the top panels. The time axis has different scales in the left and the right panels. The vertical dashed line at $t=0.06$ ms in the left panel separates the ‘peak’ and the ‘tail’ of the sub-burst $A$, as discussed in Section 4. See Figure 1 for a description of the symbols.
@@ 4 FRB 20210912A
Figure 6: Temporal evolution of the polarization state across the ‘peak’ of sub-burst $A$ of FRB 20210912A on the Poincaré sphere after correcting for $\rm RM=4.55\;rad\>m^{2}$ shown in a 3-D plot (left) and in gnomonic projection (right). Each data point represents the polarization state in a 3.8 $\mu$s time bin. The best-fit great circle is shown as a solid magenta curve in the left panel and a dashed black line in the right panel.
FRB 20210912A is another bright FRB featuring a strong primary sub-burst ($A$) followed by a relatively weaker secondary sub-burst ($B$), as shown in Figure 5. We refer to Bera et al. (2024) for dynamic spectra and other details of this FRB. The average RM of FRB 20210912A, measured over the entire burst profile, is $\rm RM_{avg}=4.55\pm 0.49\;rad\>m^{2}$. However, as discussed in Bera et al. (2024), the apparent RM varies across the two sub-bursts (see the top panels in Figure 5). We de-rotated the $Q-U$ dynamic spectra of FRB 20210912A for $\rm RM_{avg}=4.55\;rad\>m^{2}$ following the method described in Section 2.1. The frequency-averaged time profile shows high degree of polarization and both sub-bursts exhibit time-dependent gradual conversion between linear and circular polarization (middle panels in Figure 5). The PA ($\chi$) and the EA ($\psi$) show smooth correlated temporal evolution across both sub-bursts, although the nature of correlation is different in the two sub-bursts (bottom panels in Figure 5).
As noted by Bera et al. (2024), the fading ‘tail’ of sub-burst $A$ consists of a trailing component with different spectral and polarization properties compared to the ‘peak’ of the sub-burst. The PA and EA trends can be seen to change after the trailing component becomes dominant in the time profile. We hence investigate the polarization properties of the ‘peak’ and ‘tail’ of sub-burst $A$ separately, using a boundary at $t=0.06$ ms (with respect to the maximum of the total intensity profile) that roughly separates the trailing component from the brightest component. We note that sub-burst $B$ also exhibits changes in PA and EA trends at the tail; however, its relative faintness does not allow us to identify any clear component separation [^4] We note that in case of FRB 20230708A, the slight deviation of the polarization state from the great circle trajectory at the fading tail of the primary sub-burst — shown in Fig. 2 — could also be caused by the presence of a faint emission component with different polarization properties..
@@ 4.1 Polarization State in the Two Sub-bursts
We used a time resolution of 3.8 $\mu$s to probe the temporal evolution of the polarization state in the peak of sub-burst $A$ of FRB 20210912A. Like the primary sub-burst of FRB 20230708A, the frequency-averaged trajectory of the polarization state of the peak of sub-burst $A$, plotted on the surface of the Poincaré sphere following the method described in Section 2.2, is found to be well described by a great circle, as shown in Figure 6. The best-fit great circle has an inclination of $\Delta_{0}^{A}=62.4\pm 1.4^{\circ}$ and a reference PA of $\chi_{0}^{A}=-10.4\pm 1.3^{\circ}$.
@@
Figure 7: Temporal evolution of the polarization state across the ‘tail’ of sub-burst $A$ [A, left] and across sub-burst $B$ [B, right] of FRB 20210912A on the Poincaré sphere after correcting for $\rm RM=4.55\;rad\>m^{2}$. Each data point represents the polarization state of emission in a 7.6 $\mu$s time bin in the left panel and 46 $\mu$s time bin in the right panel. The best-fit great circle is shown as a dashed black line in each panel.
@@ 5 Possible Interpretations
The observed trajectories of the polarization state in FRB 20230708A and FRB 20210912A — which results from correlated variations of linear and circular polarization — are unlikely to be a random coincidence, although a formal estimation of the probability is not straightforward and critically depends on the assumptions made. Instrumental effects also appear unlikely to lead to the observed trajectories of the polarization vector as discussed in detail in Section 7.1.
The origin of the circular polarization in FRBs — whether it is an intrinsic property of radio emission or an effect of propagation through magneto-ionic media (e.g. McKinnon, 2024; Kennett & Melrose, 1998; Melrose & Luo, 2004; Lyutikov, 2022; Macquart & Melrose, 2000a, b) — remains unclear. Continuous variation of the polarization state across the burst, with changing fractions of linear and circular polarization and the polarization vector tracing smooth curves on the Poincaré sphere, has been previously observed for other FRBs (Sherman et al., 2023; Cho et al., 2020). Mandlik (2024) reported observation of FRB 20221128A with the UTMOST-NS facility at the Molonglo radio telescope, which shows the polarization vector tracing a segment of a great circle on the Poincaré sphere as seen in FRB 20230708A and FRB 20210912A. Circular polarization induced by scintillation in a turbulent magnetized medium (Macquart & Melrose, 2000b) is unlikely to lead to such smooth trajectories on the Poincaré sphere. ‘Generalized’ Faraday rotation (or Faraday conversion) has been suggested as a possible origin of varying degree of circular polarization (Cho et al., 2020; Kumar et al., 2022; Uttarkar et al., 2024). However, as discussed by Dial et al. (2025), the time-averaged polarization spectra of FRB 20230708A do not show any conclusive evidence of ‘generalized’ Faraday rotation (Lower, 2021; Uttarkar et al., 2024). Here we discuss three possible scenarios that may lead to the observed great circle trajectories of the polarization vector on the Poincaré sphere for FRB 20230708A and FRB 20210912A — transition between two orthogonal polarization modes and the effects of magneto-ionic birefringence.
@@ 5.2 Effects of Elliptical Birefringence
FRB progenitors are thought to be surrounded by a magneto-ionic shell/nebula, such as a supernova remnant or pulsar wind nebula (e.g. Chatterjee et al., 2017; Michilli et al., 2018; Bruni et al., 2024). Vedantham & Ravi (2019) and Gruzinov & Levin (2019) have suggested that linearly polarized FRB emission can be (partially) converted to circular polarization as the burst propagates through the nebula. These polarization models generally assume that the plasma in the nebula is comprised of a mixture of thermal and relativistic particles, such that the plasma’s natural modes of wave propagation are elliptically polarized. The ellipticity of the natural modes is determined by the ratio of the number densities of the thermal and relativistic particles in the plasma (Kennett & Melrose, 1998). The ellipticity is $\epsilon=\tan(\psi_{o})$, where $\psi_{o}$ is the ellipticity angle of the mode polarization vectors. If the modes are linearly polarized, $\psi_{o}=0$, $\epsilon=0$, and the plasma particles are relativistic. If the modes are circularly polarized, $\psi_{o}=\pm\pi/4$, $\epsilon=\pm 1$, and the plasma particles are thermal, as in the interstellar medium. An elliptically birefringent medium alters the phase offset, $\eta$, between its elliptically polarized natural modes of propagation and generally causes the trajectory of the observed polarization vector to trace a small circle on the Poincaré sphere (see, e.g., Figure 3 of Kennett & Melrose, 1998). The mode polarization vectors form the normal to the plane defined by the small or great circle, and the motion along the circle is associated with temporal variation of the phase offset, $\eta$. The relationship between $\psi_{o}$ and the observed inclination angles of the great circles (GCs), $\Delta$, is therefore $\psi_{o}=(\pi/2-\Delta)/2$. The ellipticity determined from the inclination angles observed in FRB 20230708A and the primary sub-burst of FRB 20210912A is $\epsilon\approx 0.25$, indicating the plasma particles are primarily relativistic.
@@
The observed GCs constrain the geometry of the mode polarization vectors relative to the polarization vector of the radiation incident upon the plasma. For the vector to trace a GC, the polarization vector of the incident radiation must be perpendicular to the mode polarization vectors. If the incident radiation is linearly polarized, its polarization vector is further constrained to reside along the line formed by the intersection of the GC with the equatorial plane of the sphere.